A COMBINATORIAL CONJECTURE RELATED WITH COMPLEX BOUNDED SYMMETRIC DOMAINS - ERRATUM

GUY ROOS

Erratum For type III_{n}, the correct value for the critical exponent is

μ_{0} = \frac{g}{d + 1} = \frac{n + 1}{\frac{n(n+1)}{2} + 1} = \frac{2}{n + \frac{2}{n+1}} ,

instead of \frac{2}{n+ \frac{1}{n+1}} .

Nevski prospekt 113/4-53, 191024 St Petersburg, Russian Federation

E-mail address: guy.roos@normalesup.org

Date: 23rd December 2004.